# Beamforming and Waveform Optimization for RF Wireless Power Transfer with Beyond Diagonal Reconfigurable Intelligent Surfaces

### v1

@@ II System Model & Problem Formulation
We consider a SISO WPT system with $N$ sub-carriers aided by a BD-RIS consisting of $M$ elements. The adoption of multitone waveforms is motivated by their ability to exploit the rectifier’s non-linearity to improve the performance [8]. For simplicity, we assume perfect channel state information at the ET as in [34, 15, 22]. Moreover, we assume there is no direct path between the ET and the ER, which is motivated by typical RIS applications to improve coverage in areas with blockages. The system model and the path of the incident and reflecting channels at the $n$th sub-carrier are illustrated in Fig. 1.a. This section describes the transmit and receive signals, BD-RIS, rectenna, and the problem formulation.
Figure 1: Illustration of (a) the system model comprising a single-antenna ET, a BD-RIS, and a single-antenna ER; and (b) examples of 4-port impedance networks for fully-connected BD-RIS (bottom-left) and D-RIS (bottom-right).
@@ II-A Transmit/Receive Signals and BD-RIS
We consider $N$ sub-carriers with $f_{n}=f_{c}+(n-1)\Delta f$ being the frequency of the $n$th sub-carrier, where $f_{c}$ is the lowest frequency and $\Delta f$ is the sub-carrier spacing. Thus, the transmit signal at time $t$ can be written as $\Re\bigl{\{}\sum_{n=1}^{N}s_{n}e^{j2\pi f_{n}t}\bigr{\}}$, where $s_{n}$ is the complex weight of the $n$th sub-carrier. [^2] We assume a linear model for the HPA being mainly interested in the impact of non-linear rectennas. Notably, this is a valid assumption for a proper transmit power such that the HPA does not enter the saturation region [27]. The transmit signal propagates through the wireless channel. We denote by $\mathbf{h}_{I,n}\in\mathbb{C}^{M\times 1}$ the channel between the transmitter and RIS and by $\mathbf{h}_{R,n}\in\mathbb{C}^{M\times 1}$ the channel between the RIS and receiver for the $n$th sub-carrier.
Assuming a multiport network model, each D-RIS element is modeled as a port connected to an independent tunable impedance [16, 40, 41]. This leads to the so-called D-RIS model, which is mathematically characterized by diagonal phase-shift matrices. In BD-RIS, the matrix of tunable elements $\mathbf{Z}\in\mathbb{C}^{M\times M}$ has non-zero off-diagonal entries [17]. Thus, the scattering matrix can be formulated as [40]
@@
where $Z_{0}$ denotes the reference impedance. In the case of reciprocal and lossless BD-RIS, [^3] For simplicity and since the goal is to investigate the gains of BD-RIS for WPT, we consider the case with no mutual coupling between different ports. the scattering matrix is symmetric ($\mathbf{\Theta}=\mathbf{\Theta}^{T})$ and unitary ($\mathbf{\Theta}^{H}\mathbf{\Theta}=\mathbf{I}_{M}$). Note that in the case of fully-connected BD-RIS, all the elements in $\mathbf{Z}$ can be non-zero, while for other sub-architectures, some off-diagonal entries become zero, and for D-RIS, $\mathbf{Z}$ becomes diagonal. Fig. 1.b illustrates an example of a 4-port impedance network for fully-connected BD-RIS and D-RIS.
The cascade channel at the $n$th sub-carrier is written as $h_{n}=\mathbf{h}_{R,n}^{T}\mathbf{\Theta}{\mathbf{h}_{I,n}}$, [^4] The underlying general model and assumptions can be seen in [41]. and the received signal at time $t$ is given by $y(t)=\sum_{n=1}^{N}\Re\bigl{\{}s_{n}h_{n}e^{j2\pi f_{n}t}\bigr{\}}$.
@@ II-B Rectenna
The ER model, i.e., rectenna, consists of the antenna equivalent circuit and a single-diode rectifier for transforming the RF received signal into DC [31, 8, 22]. By leveraging this model, assuming perfect matching, and using the Taylor’s expansion, the output current of the rectenna is given by
@@
where $K_{2}$, $K_{4}$ are determined by the characteristics of the rectifier circuit. One can consider $\bar{n}=4$ and model the main sources of non-linearity using the fourth order term, while the second order term represents the ideal linear rectenna model [22, 34]. Note that (3) requires averaging over time samples, which makes the optimization extremely challenging due to the need to consider a high sampling rate for achieving sufficient accuracy. Fortunately, by expressing (3) in the frequency domain and assuming $n_{0}=4$, we obtain [22]
$i_{dc}$ $=\frac{K_{2}}{2}\sum_{n}\left\lVert s_{n}h_{n}\right\rVert^{2}+\frac{3K_{4}}{8}\sum_{\begin{subarray}{c}n_{0},n_{1},n_{2},n_{3}\\ n_{0}+n_{1}=n_{2}+n_{3}\end{subarray}}{({h}_{n_{0}}{s}_{n_{0}})}^{\star}\ldots$
@@
which is more tractable than the sampling-dependent model.
@@ II-C Problem Formulation
We aim to maximize the harvested power given a transmit power budget, $P_{T}$. By leveraging the fact that the DC harvested power is an increasing function of the DC current at the ER, the problem can be formulated as
@@
$\frac{1}{2}\sum_{n=1}^{N}\left\lVert s_{n}\right\rVert^{2}\leq P_{T}.$ (5d)
@@ III-A Waveform Optimization with Fixed $h_{n}$
First, we assume the cascade channel $h_{n}$ to be fixed and optimize the signal weights $s_{n},\forall n$. We proceed by defining $s_{n}=\bar{s}_{n}e^{j\tilde{s}_{n}}$, where $\bar{s}_{n}$ and $\tilde{s}_{n}$ are the amplitude and the phase of $s_{n}$, respectively. Similarly, we can write $h_{n}=\bar{h}_{n}e^{j\tilde{h}_{n}}$. It is evident that the optimal $\tilde{s}_{n}$ must compensate for the phases in (II-B), leading to a real-valued $i_{dc}$ [7]. Thus, it is sufficient to have $\tilde{s}_{n}^{*}=-\tilde{h}_{n},\forall n$ leading to $\bar{s}_{n}e^{j\tilde{s}_{n}^{*}}h_{n}=\bar{s}_{n}\bar{h}_{n},\forall n$. Now the only goal of the optimization is to find the optimal amplitudes for the signal weights at different sub-carriers. For this, the problem for a given $\boldsymbol{\Theta}$ can be reformulated as
@@
Note that ${y(t)}$ is linear w.r.t. $\bar{s}_{n}$. By leveraging the second-order convexity condition [42] and the linearity of the mathematical average operator, it can be easily verified that $i_{dc}$ is convex w.r.t. $y(t)$. Then, $i_{dc}$ is convex w.r.t. $\bar{s}_{n}$ since the composition of an affine with a convex function is convex. ∎
@@
Finally, the problem can be iteratively solved using standard convex optimization tools, e.g., CVX [43].
However, utilizing standard solvers for the optimization problem in (9) requires solving a quadratic program [42] in each SCA iteration. Notably, the complexity of quadratic programs scales with a polynomial function of the problem size, while the degree of the polynomial mainly depends on the type of solver. Let us consider a simple solver based on the Newton method, which has $\mathcal{O}(n^{3})$ complexity [42], where $n$ is the problem size, leading to $n$ scaling with the number of variables $N$. Therefore, solving problems such as (9) using standard tools is not computationally efficient.
@@
The local solution of (9) can be updated using
$\bar{s}_{n}^{(l+1)}=\bar{s}_{n}^{(l)}+\rho_{s}(\bar{s}_{n}^{*}-\bar{s}_{n}^{(l)}),$ (10)
where $\rho_{s}\leq 1$ is a convergence control parameter,
$\bar{s}_{n}^{*}$ $=\frac{g(\bar{s}_{n}^{(l)})}{\lambda^{(l)}},\quad\forall n,$ (11)
$\lambda^{(l)}$ $=\sqrt{\frac{1}{2P_{T}}\sum_{n=1}^{N}{g(\bar{s}_{n}^{(l)})}^{2}}.$ (12)
@@
    The proof is provided in Appendix A.
Leveraging Theorem 1 to iteratively solve (9) leads to considerably lower computation complexity since closed-form expressions are utilized in each SCA iteration. Algorithm 1 illustrates the proposed iterative method for obtaining $s_{n},\forall n$. First, the scaled matched filter approach, proposed in [23], is used to initialize the signal weights such that
Then, the optimization variables and the corresponding dual variable are updated iteratively until convergence. By leveraging the lower-bound properties of Taylor’s approximation and writing the KKT conditions of (6), it can be seen that (9a) is monotonically increasing and the obtained solution by Algorithm 1 satisfies the KKT conditions of (6).
@@ III-B SDR-based Beamforming for Fully Connected BD-RIS
$\operatorname*{maximize}_{\boldsymbol{\Theta}}\quad$ $i_{dc}$ (14a)
@@
$[\mathbf{P}]_{M(m-1)+n,k}=\begin{cases}1,&k={m(m-1)}/{2}+n,\ 1\leq n\leq m\\ 1,&k={n(n-1)}/{2}+m,\ m<n\leq M\\ 0,&\text{otherwise}.\end{cases}$ (16)
  • Proof.
    The proof is provided in Appendix B
Leveraging Proposition 1 allows us to remove the constraint (5b) by using $\boldsymbol{\theta}$ as the optimization variable. However, the complexity caused by the unitary constraint still remains. For this, we leverage the idea in [34] and define $\mathbf{z}_{n}=s_{n}\mathbf{a}_{n}$, $\mathbf{D}_{0}=\mathbf{z}_{1}\mathbf{z}_{1}^{H}+\ldots+\mathbf{z}_{N}\mathbf{z}_{N}^{H}$, $\mathbf{D}_{1}=\mathbf{z}_{1}\mathbf{z}_{2}^{H}+\ldots+\mathbf{z}_{N-1}\mathbf{z}_{N}^{H}$, and $\mathbf{D}_{N-1}=\mathbf{z}_{1}\mathbf{z}_{N}^{H}$. Hereby, $i_{dc}$ can be reformulated as
@@
where $\bar{\mathbf{P}}_{i,j}=\mathbf{P}_{i}^{T}\mathbf{P}_{j}$ and $\mathbf{P}_{i}$ is a permutation matrix containing the $(iM-M+1)$th to the $iM$th row of $\mathbf{P}$.
Now, we define $d_{n}=\boldsymbol{\theta}^{H}\mathbf{D}_{n}\boldsymbol{\theta}$, $\mathbf{d}=[d_{1},\ldots,d_{N}]^{T}$, and positive semidefinite matrices $\mathbf{K}_{0}=\mathrm{diag}\{\frac{3}{8}K_{4},\frac{3}{4}K_{4},\ldots,\frac{3}{4}K_{4}\}\succeq 0$ and $\mathbf{X}=\boldsymbol{\theta}\boldsymbol{\theta}^{H}$. Hereby, the problem can be reformulated as
@@
Note that problem (19) is still non-convex and challenging to solve since it deals with the maximization of a convex objective function and includes a rank-1 constraint. For (19a), SCA can be used to iteratively update the objective function by approximating it using its first-order Taylor’s expansion. Specifically, the quadratic term in (19a) can be approximated in the neighborhood of $\mathbf{d}^{(l)}$ by [34]
@@
Since $f(\mathbf{d},\mathbf{d}^{(l)})\leq\mathbf{d}^{H}\mathbf{K}_{0}\mathbf{d}$ always holds, it can be used as a lower bound and maximizing the Taylor’s approximation iteratively leads to maximizing the original quadratic term. Note that (19e) can be equivalently written as
@@
which is non-convex since it deals with the difference of two convex functions. However, the right-hand-side term can be approximated using its first-order Taylor’s expansion at a local point $\bar{\mathbf{X}}$. Specifically, $\left\lVert\mathbf{X}\right\rVert_{F}$ can be replaced by a convex lower-bound given by [39]
@@
This problem is a standard SDP that can be solved using convex optimization tools such as CVX [43]. Note that iteratively solving (23) leads to convergence to a rank-1 solution, which is a local optimum solution of (19). [^5] The proof is similar to the one provided in [34]. However, this approach requires many iterations for convergence since $\mathbf{X}$ is a large matrix and the approximation is done in both the objective function and a constraint. To cope with this complexity, we use SDR to reformulate the problem in the neighborhood of the initial point $\mathbf{d}^{(l)}$ as
@@
which is also a standard SDP similar to (23). Moreover, if the obtained $\mathbf{X}^{*}$ is a rank-1 matrix, the SDR is tight and $\mathbf{X}^{*}$ is a stationary point of problem (25), leading to a local optimum point extracted by $\mathbf{X}^{*}=\boldsymbol{\theta}^{*}{\boldsymbol{\theta}^{*}}^{H}$. Then, $\boldsymbol{\theta}^{*}$ can be used to obtain $\mathbf{\Theta}^{*}=\mathrm{Vec}^{-1}(\mathbf{P}\boldsymbol{\theta}^{*})$. However, it might happen that $\mathrm{rank}(\mathbf{X}^{*})>1$, which is the case that $\mathbf{X}^{*}$ satisfies the KKT conditions of problem (25) (see [34] for the proof). For this, we obtain an approximate $\boldsymbol{\theta}^{*}$ using the Gaussian randomization method in [44]. Note that in this case, using $\boldsymbol{\theta}^{*}$ to construct $\mathbf{\Theta}$ leads to a symmetric $\mathbf{\Theta}$, but not necessarily a unitary matrix. Thus, it is essential to map the final solution into the feasible space of the problem. Let us proceed by writing $\mathbf{\Theta}^{\prime}=\mathrm{Vec}^{-1}(\mathbf{P}\boldsymbol{\theta}^{*})$. Then, by leveraging the fact that $\mathbf{\Theta}^{\prime}$ is symmetric, we can write the singular value decomposition (SVD) as $\mathbf{\Theta}^{\prime}=\mathbf{Q}\boldsymbol{\Sigma}\mathbf{Q}^{T}$, where $\mathbf{Q}$ is a unitary matrix and $\boldsymbol{\Sigma}$ is a diagonal matrix containing the singular values of $\mathbf{\Theta}^{\prime}$. It is evident that if the diagonal elements of $\boldsymbol{\Sigma}$ are unit modulus, $\mathbf{\Theta}^{\prime}$ is unitary. However, this is only the case when $\mathrm{rank}(\mathbf{X}^{*})=1$ and SDR is tight, while for higher-rank cases, we propose a randomization-based method to obtain a feasible solution.
@@
3:$f^{\star}=0$, $i_{dc}=\infty$ , $[\boldsymbol{Z}_{i,m}]=jZ_{0},\forall i,m$,
4:$\mathbf{\Theta}^{(l)}=(\mathbf{Z}+Z_{0}\mathbf{I}_{M})^{-1}(\mathbf{Z}-Z_{0}\mathbf{I}_{M})$

### v2

@@ II System Model & Problem Formulation
We consider a SISO WPT system with $N$ sub-carriers aided by a BD-RIS consisting of $M$ elements. The adoption of multi-tone waveforms is motivated by their ability to exploit the rectifier’s non-linearity to improve the performance [clerckx2018beneficial]. For simplicity, we assume perfect channel state information at the ET as in [clerckxRISWPT, bdris_main_ref, clreckxWFdesign]. Moreover, we assume there is no direct path between the ET and the ER, which is motivated by typical RIS applications to improve coverage in areas with blockages. By excluding the direct link, we isolate and highlight the contribution of the BD-RIS to the overall performance, allowing for a fair and focused comparison between different RIS architectures. The system model and the path of the incident and reflecting channels at the $n$th sub-carrier are illustrated in Fig. 1.a. This section describes the transmit and receive signals, BD-RIS, rectenna, and the problem formulation.
Remark 1.
The SISO setup is chosen to focus on the core contributions of this work, namely, the investigation of frequency diversity, BD-RIS beamforming gains, and nonlinear energy harvesting. This provides a clear and tractable framework for analyzing the behavior of BD-RIS systems and enables fair comparisons with conventional RIS designs. Importantly, the use of BD-RIS introduces significant mathematical complexity due to its general scattering structure, particularly when combined with nonlinear rectenna models and multi-carrier waveform optimization. Therefore, the considered SISO scenario serves as a necessary starting point for characterizing the fundamental gains of BD-RIS in WPT.
Figure 1: Illustration of (a) the system model comprising a single-antenna ET, a BD-RIS, and a single-antenna ER; and (b) examples of 4-port impedance networks for fully connected BD-RIS (bottom-left) and D-RIS (bottom-right).
@@ II-A Transmit/Receive Signals and BD-RIS
We consider $N$ sub-carriers with $f_{n}=f_{c}+(n-1)\Delta f$ being the frequency of the $n$th sub-carrier, where $f_{c}$ is the lowest frequency and $\Delta f$ is the sub-carrier spacing. Thus, the transmit signal at time $t$ can be written as $\Re\bigl\{\sum_{n=1}^{N}s_{n}e^{j2\pi f_{n}t}\bigr\}$, where $s_{n}$ is the complex weight of the $n$th sub-carrier. [^2] We assume a linear model for the power amplifier, being mainly interested in the impact of non-linear rectennas. Notably, this is a valid assumption for a proper transmit power such that the power amplifier does not enter the saturation region [SISOAllClreckx]. The transmit signal propagates through the wireless channel. We denote by $\mathbf{h}_{I,n}\in\mathbb{C}^{M\times 1}$ the channel between the transmitter and RIS and by $\mathbf{h}_{R,n}\in\mathbb{C}^{M\times 1}$ the channel between the RIS and receiver for the $n$th sub-carrier.
Assuming a multiport network model, each D-RIS element is modeled as a port connected to an independent tunable impedance [BD-RIS_scattering_clerckx, univnerini, bdrissportdesign]. This leads to the so-called D-RIS model, which is mathematically characterized by diagonal phase-shift matrices. In BD-RIS, the matrix of tunable elements $\mathbf{Z}\in\mathbb{C}^{M\times M}$ has non-zero off-diagonal entries [BD-RIS_General]. Thus, the scattering matrix can be formulated as [univnerini]
@@
where $Z_{0}$ denotes the reference impedance. In the case of reciprocal and lossless BD-RIS, [^3] For simplicity and since the goal is to investigate the gains of BD-RIS for WPT, we consider the case with no mutual coupling between different ports. the scattering matrix is symmetric ($\mathbf{\Theta}=\mathbf{\Theta}^{T})$ and unitary ($\mathbf{\Theta}^{H}\mathbf{\Theta}=\mathbf{I}_{M}$). Note that in the case of fully connected BD-RIS, all the elements in $\mathbf{Z}$ can be non-zero, while for other sub-architectures, some off-diagonal entries become zero, and for D-RIS, $\mathbf{Z}$ becomes diagonal. Fig. 1.b illustrates an example of a 4-port impedance network for fully connected BD-RIS and D-RIS.
The cascade channel at the $n$th sub-carrier is written as $h_{n}=\mathbf{h}_{R,n}^{T}\mathbf{\Theta}{\mathbf{h}_{I,n}}$, [^4] The underlying general model and assumptions can be seen in [bdrissportdesign]. and the received signal at time $t$ is given by $y(t)=\sum_{n=1}^{N}\Re\bigl\{s_{n}h_{n}e^{j2\pi f_{n}t}\bigr\}$.
@@ II-B Rectenna
The ER model, i.e., rectenna, consists of the antenna equivalent circuit and a single-diode rectifier for transforming the RF received signal into DC [azarbahram2024waveform, clerckx2018beneficial, clreckxWFdesign]. By leveraging this model, assuming perfect matching, and using Taylor’s expansion, the output current of the rectenna is given by
@@
where $K_{2}$, $K_{4}$ are determined by the characteristics of the rectifier circuit. One can consider $\bar{n}=4$ and model the main sources of non-linearity using the fourth order term, while the second order term represents the ideal linear rectenna model [clreckxWFdesign, clerckxRISWPT]. Note that (3) requires averaging over time samples, which makes the optimization extremely challenging due to the need to consider a high sampling rate for achieving sufficient accuracy. Fortunately, by expressing (3) in the frequency domain and assuming $\bar{n}=4$, we obtain [clreckxWFdesign]
$i_{dc}$ $=\frac{K_{2}}{2}\sum_{n}|{s_{n}h_{n}}|^{2}+\frac{3K_{4}}{8}\sum_{\begin{subarray}{c}n_{0},n_{1},n_{2},n_{3}\\ n_{0}+n_{1}=n_{2}+n_{3}\end{subarray}}{({h}_{n_{0}}{s}_{n_{0}})}^{\star}\ldots$
@@
which is more tractable than the sampling-dependent model. Notice that $i_{dc}$ is a real-valued scalar obtained by averaging the second and fourth moments of the received real-valued signal.
@@ II-C Problem Formulation
We aim to maximize the harvested power given a transmit power budget, $P_{T}$, and assuming perfect channel state information (CSI). By leveraging the fact that the DC harvested power is an increasing function of the DC at the ER, the problem can be formulated as
@@
$\frac{1}{2}\sum_{n=1}^{N}|{s_{n}}|^{2}\leq P_{T}.$ (5d)
@@
Remark 2.
Note that the power consumption of BD-RIS scales with the number of interconnections between its ports, which directly increases the circuit complexity and, consequently, the overall power consumption compared to traditional D-RIS architectures [RIS_power_conusmption_TWC_2024]. However, this work focuses solely on the output power of the active transmitter, and thus, we assume that the BD-RIS has sufficient power available for its internal operation and impedance network tuning.
Remark 3.
For conventional diagonal RIS, two main CSI acquisition strategies exist [CSI_survey_RIS]: i) channel estimation using partially activated RIS elements by RF chains, offering low overhead and compatibility with BD-RIS, but incurring additional hardware costs; ii) Cascaded channel estimation via RIS using pilot designs, which avoids RF chains but needs to be adapted for BD-RIS due to its tightly coupled channel structure. A BD-RIS-specific approach is explored in [channel_estim_BDRIS].
@@ III-A Waveform Optimization with Fixed $h_{n}$
First, we assume the cascade channel $h_{n}$ to be fixed and optimize the signal weights $s_{n},\forall n$. We proceed by defining $s_{n}=\bar{s}_{n}e^{j\tilde{s}_{n}}$, where $\bar{s}_{n}$ and $\tilde{s}_{n}$ are the amplitude and the phase of $s_{n}$, respectively. Similarly, we can write $h_{n}=\bar{h}_{n}e^{j\tilde{h}_{n}}$. It is evident that the optimal $\tilde{s}_{n}$ must compensate for the phases in (II-B), leading to a real-valued $i_{dc}$ [clerckx-nonlinearharvester]. Thus, it is sufficient to have $\tilde{s}_{n}^{*}=-\tilde{h}_{n},\forall n$ leading to $\bar{s}_{n}e^{j\tilde{s}_{n}^{*}}h_{n}=\bar{s}_{n}\bar{h}_{n},\forall n$. Now the only goal of the optimization is to find the optimal amplitudes for the signal weights at different sub-carriers. For this, the problem for a given $\boldsymbol{\Theta}$ can be reformulated as
@@
Note that y(t){y(t)} is linear w.r.t. $\bar{s}_{n}$. By leveraging the second-order convexity condition [boyd2004convex] and the linearity of the mathematical average operator, it can be easily verified that $i_{dc}$ is convex w.r.t. $y(t)$. Then, $i_{dc}$ is convex w.r.t. $\bar{s}_{n}$ since the composition of an affine with a convex function is convex. ∎
@@
Finally, the problem can be iteratively solved using standard convex optimization tools, e.g., CVX [cvxref].
However, utilizing standard solvers for the optimization problem in (9) requires solving a quadratic program [boyd2004convex] in each SCA iteration. Notably, the complexity of quadratic programs scales with a polynomial function of the problem size, while the degree of the polynomial mainly depends on the type of solver. Let us consider a simple solver based on the Newton method, which has $\mathcal{O}(n^{3})$ complexity [boyd2004convex], where $n$ is the problem size, leading to $n$ scaling with the number of variables $N$. Therefore, solving problems such as (9) using standard tools is not computationally efficient.
Let us proceed by writing the Lagrangian of (9) as
By setting (11) to zero, we attain
$\bar{s}_{n}^{*}=\frac{g(\bar{s}_{n}^{(l)})}{\lambda^{(l)}},\quad\forall n.$ (12)
Moreover, the optimal solution must utilize the whole transmit power budget, i.e., $\frac{1}{2}\sum_{n=1}^{N}{\bar{s}_{n}}^{2}=P_{T}$. Thus, by substituting (12) into the latter equality, the dual variable $\lambda^{(l)}$ is given by
$\lambda^{(l)}=\sqrt{\frac{1}{2P_{T}}\sum_{n=1}^{N}{g(\bar{s}_{n}^{(l)})}^{2}}.$ (13)
Finally, we introduce an additional parameter $\rho_{s}\leq 1$ to control the convergence of the solution such that the sub-gradient update is given by
$\bar{s}_{n}^{(l+1)}=\bar{s}_{n}^{(l)}+\rho_{s}(\bar{s}_{n}^{*}-\bar{s}_{n}^{(l)}),$ (14)
which can be used to obtain the local solution of (9). Hereby, we can iteratively solve (6), which leads to considerably lower computation complexity since closed-form expressions are utilized in each SCA iteration. Algorithm 1 illustrates the proposed iterative method for obtaining $s_{n},\forall n$. First, the scaled matched filter approach, proposed in [brunolowcomp], is used to initialize the signal weights such that
Then, the optimization variables and the corresponding dual variable are updated iteratively until convergence.
@@
Algorithm 1 converges to a stationary point of problem (6).
@@
The first-order Taylor approximation is a global under-estimator of a convex function, leading to $\tilde{i}_{dc}(\bar{s}_{n},\bar{s}_{n}^{(l)})\leq i_{dc}(\bar{s}_{n})$. By denoting $\bar{s}_{n}^{*}$ as the solution of problem (9) at the lth iteration, we can write $\tilde{i}_{dc}(\bar{s}_{n}^{*},\bar{s}_{n}^{(l)})\geq\tilde{i}_{dc}(\bar{s}_{n}^{(l)},\bar{s}_{n}^{(l)})=i_{dc}(\bar{s}_{n}^{(l)})$. Combining this with the fact that $i_{dc}$ is convex and $\rho_{s}\leq 1$ in (14), we obtain that the sequence $\{i_{dc}(\bar{s}_{n}^{(l)})\}$ is monotonically increasing. Since $i_{dc}$ is continuous and bounded due to constraint (6b), $\lim_{l\rightarrow\infty}i_{dc}(\bar{s}_{n}^{(l)})=i_{dc}^{*}$ exists. Therefore, the solution converges to a point that fulfills the KKT conditions of the original problem (6), indicating that it is a stationary point of (6). ∎
@@ III-B SDR-based Beamforming for Fully Connected BD-RIS
$\operatorname*{maximize}_{\boldsymbol{\Theta}}\quad$ $i_{dc}$ (16)
@@
[…]
The proof is provided in Appendix A
Leveraging Proposition 1 allows us to remove the constraint (5b) by using $\boldsymbol{\theta}$ as the optimization variable. However, the complexity caused by the unitary constraint still remains. For this, we leverage the idea in [clerckxRISWPT] and define $\mathbf{z}_{n}=s_{n}\mathbf{a}_{n}$, $\mathbf{D}_{0}=\mathbf{z}_{1}\mathbf{z}_{1}^{H}+\ldots+\mathbf{z}_{N}\mathbf{z}_{N}^{H}$, $\mathbf{D}_{1}=\mathbf{z}_{1}\mathbf{z}_{2}^{H}+\ldots+\mathbf{z}_{N-1}\mathbf{z}_{N}^{H}$, and $\mathbf{D}_{N-1}=\mathbf{z}_{1}\mathbf{z}_{N}^{H}$. Hereby, $i_{dc}$ can be reformulated as
@@
where $\bar{\mathbf{P}}_{i,j}=\mathbf{P}_{i}^{T}\mathbf{P}_{j}$ and $\mathbf{P}_{i}$ is a permutation matrix containing the (iMM+1)(iM-M+1)th to the $iM$th row of $\mathbf{P}$.
[…]
@@
Note that problem (21) is still non-convex and challenging to solve since it deals with the maximization of a convex objective function and includes a rank-1 constraint. For (21a), SCA can be used to iteratively update the objective function by approximating it using its first-order Taylor expansion. Specifically, the quadratic term in (21a) can be approximated in the neighborhood of $\mathbf{d}^{(l)}$ by [clerckxRISWPT]
@@
Since $f(\mathbf{d},\mathbf{d}^{(l)})\leq\mathbf{d}^{H}\mathbf{K}_{0}\mathbf{d}$ always holds, it can be used as a lower bound, and maximizing Taylor’s approximation iteratively leads to maximizing the original quadratic term. Note that (21e) can be equivalently written as
@@
which is non-convex since it deals with the difference of two convex functions. However, the right-hand-side term can be approximated using its first-order Taylor expansion at a local point $\bar{\mathbf{X}}$. Specifically, $\left\lVert\mathbf{X}\right\rVert_{F}$ can be replaced by a convex lower-bound given by [shabir2024electromagneticallyconsistentoptimizationalgorithms]
@@
This problem is a standard SDP that can be solved using convex optimization tools such as CVX [cvxref]. However, this approach requires many iterations for convergence since $\mathbf{X}$ is a large matrix and the approximation is done in both the objective function and a constraint. To cope with this complexity, we use SDR to reformulate the problem in the neighborhood of the initial point $\mathbf{d}^{(l)}$ as
@@
which is also a standard SDP similar to (25). Moreover, if the obtained $\mathbf{X}^{*}$ is a rank-1 matrix, the SDR is tight and $\mathbf{X}^{*}$ is a stationary point of the original convex subproblem including rank-1 constraint, leading to the solution extracted by $\mathbf{X}^{*}=\boldsymbol{\theta}^{*}{\boldsymbol{\theta}^{*}}^{H}$. Then, $\boldsymbol{\theta}^{*}$ can be used to obtain $\mathbf{\Theta}^{*}=\mathrm{Vec}^{-1}(\mathbf{P}\boldsymbol{\theta}^{*})$. However, it might happen that $\mathrm{rank}(\mathbf{X}^{*})>1$, which is the case that $\mathbf{X}^{*}$ is a stationary point of problem (27) (see [clerckxRISWPT] for the proof). For this, we obtain an approximate $\boldsymbol{\theta}^{*}$ using the Gaussian randomization method in [guss_randomization]. Note that in this case, using $\boldsymbol{\theta}^{*}$ to construct $\mathbf{\Theta}$ leads to a symmetric $\mathbf{\Theta}$, but not necessarily a unitary matrix. Thus, it is essential to map the final solution into the feasible space of the problem. Let us proceed by writing $\mathbf{\Theta}^{\prime}=\mathrm{Vec}^{-1}(\mathbf{P}\boldsymbol{\theta}^{*})$. Then, by leveraging the fact that $\mathbf{\Theta}^{\prime}$ is symmetric, we can write the singular value decomposition (SVD) as $\mathbf{\Theta}^{\prime}=\mathbf{Q}\boldsymbol{\Sigma}\mathbf{Q}^{T}$, where $\mathbf{Q}$ is a unitary matrix and $\boldsymbol{\Sigma}$ is a diagonal matrix containing the singular values of $\mathbf{\Theta}^{\prime}$. It is evident that if the diagonal elements of $\boldsymbol{\Sigma}$ are unit modulus, $\mathbf{\Theta}^{\prime}$ is unitary. However, this is only the case when $\mathrm{rank}(\mathbf{X}^{*})=1$ and SDR is tight, while for higher-rank cases, we propose a randomization-based method to obtain a feasible solution.
@@
[…]